TestingAccount:

If I have 500 miles to run and i run 90 at a speed of 20mph how many miles do i run in a minute

TestingAccount:

meters in a minue

TestingAccount:

not miles

jhonyy9:

use the rule of 3 simple and write 20 mph mean 20 miles per hour - 1 hour = 60 minutes x miles in 1 minute cross multiplie and calcule the x value x = 20*1/60 = 20/60 = 2/6 = 1/3 so in a minute he/she run 1/3 mile hope helped understandably easy @Laylalyssa

TestingAccount:

tyyyyyyyyyyyyyyyyy

TestingAccount:

tytytyty